A STUDY OF THE OVERTONE BANDS OF HALOGEN OXIDE MOLECULES BY LMR.

\underline{FILIPPO TAMASSIA} and J. M. BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, UK, OX1 3QZ.

The 2-0 overtone bands of the FO and BrO radicals in their X̃^{2}Π_{3/2} electronic ground state have been studied by CO Laser Magnetic Resonance. A number of vibrational, rotational, centrifugal distortion, spin-orbit, hyperfine, lambda-doubling and Zeeman parameters have been determined. The band origins and the anharmonicity constants ω_{e}x_{e} (cm^{−1}) are the following:

FO  ν_{2−0} = 2045.47377(27)  ω_{e}x_{e} = 10.22713(10)
79BrO  ν_{2−0} = 1437.4786(18)  ω_{e}x_{e} = 4.67583(43)
81BrO  ν_{2−0} = 1434.5422(19)  ω_{e}x_{e} = 4.65703(39)